OpenStudy (anonymous):

Solve for x. √2x+3= √6x-1 all numbers are under the rad sign

OpenStudy (agreene):

\[\sqrt{2x+3}=\sqrt{6x-1}\] \[\sqrt{2x+3}^2=\sqrt{6x-1}^2 \implies \cancel{\sqrt{}}(2x+3)\cancel{^2}=\cancel{\sqrt{}}(6x-1)\cancel{^2}\] \[2x+3=6x-1\] \[-4x=-4\] \[x=1\]

OpenStudy (anonymous):

thankyouu